OpenStudy (anonymous):

The equation y=-1777x+27,153 can be used to predict the number y of gun deaths in the United States x years after 2000, that is x=0 corresponds to 2000, x=3 corresponds to 2003.. and so on. Predict the number of gun deaths in 2005 , and 2009. In what year will the number of gun deaths be 9383?

OpenStudy (saifoo.khan):

Do you wanna try it? :(

OpenStudy (saifoo.khan):

Alright, let me do it.

OpenStudy (anonymous):

oh sorry I wasnt paying attentnion

OpenStudy (saifoo.khan):

lol, i have solve it for Billion times.

OpenStudy (saifoo.khan):

Let me do it one more times, y=-1777x+27,153 y = -1777(5) + 27,153 y = 18268 <------------- year 2005

OpenStudy (saifoo.khan):

y=-1777x+27,153 y = -1777(9) + 27,153 y = 11160 <----------- year 2009

OpenStudy (anonymous):

awe sweet,

OpenStudy (anonymous):

wait that doesnt make sense...

OpenStudy (saifoo.khan):

Last part is left. 9383 = -1777x+27,153 x = 10 So, in 2010 the gun deaths will be 9383

OpenStudy (saifoo.khan):

Why doesn't it? :O

OpenStudy (anonymous):

how si thre more deaths in 2005 than 2009...

OpenStudy (saifoo.khan):

This is because the gun death decreases, they aren't increasing.

OpenStudy (anonymous):

oh okay ...

OpenStudy (saifoo.khan):

:D